### v1

# Language model developers should report train-test overlap

Abstract
Language models are extensively evaluated, but correctly interpreting evaluation results requires knowledge of train-test overlap, which refers to the extent to which the language model is trained on the very data it is being tested on. The public currently lacks adequate information about train-test overlap: most models have no public train-test overlap statistics, and third parties cannot directly measure train-test overlap since they do not have access to the training data. To make this clear, we document the practices of 30 models, finding that just 9 models report train-test overlap: 4 models release training data under open-source licenses, enabling the community to directly measure train-test overlap, and 5 models publish their train-test overlap methodology and statistics. By engaging with language model developers, we provide novel information about train-test overlap for three additional models. Overall, we take the position that language model developers should publish train-test overlap statistics and/or training data whenever they report evaluation results on public test sets. We hope our work increases transparency into train-test overlap to increase the community-wide trust in model evaluations.
1 Introduction
The artificial intelligence (AI) community has built hundreds of evaluations to better understand language models (Srivastava et al., 2023; Hendrycks et al., 2021a; Gao et al., 2024; Myrzakhan et al., 2024; Chiang et al., 2024; Rein et al., 2023; Liang et al., 2023). These evaluations cannot be correctly interpreted without knowledge of train-test overlap, which we define as the extent to which the evaluation test data appears in the training data.
Prior to the rise of language models trained on web-scale data, the AI community used standard train/test set splits, where a model would be trained on the training set and tested on the test set to ensure validity of results (Jurafsky & Martin, 2009; Russell & Norvig, 2009). In that regime, the designer of an evaluation generally would specify both the training and test sets. In contrast, today foundation model developers decide on their own training sets, which they often do not release, and evaluation designers decide on test sets, which they often release. Overall, the shift to web-scale training data with poor documentation of data provenance, along with the two-party specification of training and test data, contributes to poor understanding of train-test overlap (Longpre et al., 2023).
Train-test overlap can arise for several reasons. First, since evaluation datasets are often made public on the Internet (e.g. via public repositories like GitHub and Hugging Face), these datasets may be scraped and then trained upon. Second, since evaluation datasets often use already-public material (e.g. the held-out examples in SQuAD still depend on public Wikipedia data (Rajpurkar et al., 2016)), the underlying data may be easily trained upon. Third, since evaluation datasets are often input into models to conduct evaluations (e.g. to evaluate GPT-4 via the OpenAI API), these datasets may be stored and used to train future models. Better understanding how train-test overlap arises may facilitate solutions for appropriately navigating the challenges it presents (Oren et al., 2023).
A growing literature demonstrates high train-test overlap for language models, which contributes to significant degradation in performance between seen and unseen test examples (Lewis et al., 2020; Elangovan et al., 2021; Vu et al., 2023). For example, OpenAI initially reported that GPT-4 had achieved state-of-the-art performance on a test set of Codeforces coding questions, claiming that there was no contamination (OpenAI, 2023). Yet it was later demonstrated [^1] See https://twitter.com/cHHillee/status/1635790330854526981?lang=en that while GPT-4 achieves 100% accuracy for 10 pre-2021 problems, the model achieves 0% accuracy on more recent problems. More generally, Kapoor & Narayanan (2022) document that test data often leaks into training data across many domains. Therefore, improving the community’s understanding train-test overlap will increase the validity of, and trust in, evaluations.
Given the value of understanding train-test overlap, we study the practices of 30 language model. We find that 9 models have published sufficient data for the AI community to contextualize train-test overlap: 4 models (OLMo—AI2, GPT-NeoX—EleutherAI, RedPajama INCITE—Together, StarCoder 2—BigCode/HuggingFace/ServiceNow) have released open-source datasets that the community can inspect for train-test overlap, and 5 models (GPT-4—OpenAI, Llama 3.1—Meta, Qwen2—Alibaba, Palmyra—Writer, Apple Intelligence—Apple) have published their methodology and statistics for train-test overlap. The remaining 23 models do report evaluation results on public test sets, but do not (adequately) report train-test overlap results. In parallel to models reporting train-test overlap, the community is building black-box methods to estimate train-test overlap without access to training data (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023), but these approaches are quite limited at present. We take the position that language model developers should report train-test overlap.
Position: Language model developers should report train-test overlap. Language model developers routinely publish evaluations of their models on public test sets. However, these evaluations are often not accompanied with train-test overlap statistics, making it difficult to assess their validity. Similar to how statisticians are expected to release confidence intervals to ensure validity of their results, we argue that language model developers that publish results on public test sets should release their models’ training data and/or publish accompanying train-test overlap statistics so that the AI community can correctly interpret the evaluation results.
2 Strategies to Estimate and Address Train-Test Overlap
The prevalence of train-test overlap as an issue in the AI community [^2] Potential evidence of train-test overlap is often flagged by members of the AI community on social media. See, e.g., https://twitter.com/dhuynh95/status/1775568278557192411 has led to the development of various strategies to estimate and address train-test overlap, including black-box methods, private test sets, novel test sets, and canary strings. We cover each of these in turn then discuss our approach.
Black-box methods
involve researchers working to estimate train-test overlap through model api access and the test set rather than directly through access to the training set. Notably, there have been efforts to estimate train-test overlap via prompting, word probabilities, and test example orderings (Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023). Golchin & Surdeanu (2023) prompt the model with the dataset name, partition type, and an initial segment of the reference string, and mark train-test overlap if the model responds with a exact or similar copy in the output. Shi et al. (2023) estimate train-test overlap via the probability outputs of outlier words, with the hypothesis that unseen examples is likely to contain few outliers with low probabilities. Oren et al. (2023) estimate train-test overlap by considering the ordering of test instances, noting that language models with train-test overlap are likely to memorize such ordering. These methods can be helpful for estimation and as a sanity check to white-box approaches, but have currently have limitations as they are not robust to adversarial settings such as if a developer fine-tuned its model to avoid revealing training data and even in the benign setting, require certain assumptions such as requiring a certain threshold of frequencies for detection or certain methods of training (Casper et al., 2024; Golchin & Surdeanu, 2023; Shi et al., 2023; Oren et al., 2023). Estimating and interpreting train-test overlap is difficult even in the white-box setting with direct access to the training data as current approaches have significant limitations; with further constraints in the black-box setting, the challenges only increase.
Private test sets
such as SQuAD (Rajpurkar et al., 2018) and SEAL (Scale, 2024) allow researchers to keep a portion or all of the test set hidden, meaning that the test set is not publicly accessible on the internet and developers are therefore much less likely to train models on it. While private test sets can be valuable, they raise potential concerns regarding data transparency. For instance, unless the private test set is shared with a trustworthy third party, the community must rely upon a single organization’s assessment of the test set’s validity. In any event, public test sets are the industry standard and will continue to exist, though private and public test sets can coexist in a healthy testing ecosystem.
Novel test sets
that include data that was produced after the knowledge cutoff date of a model also help mitigate train-test overlap. Including recent data is a best practice for new test sets, though this may be difficult if, for example, a new test set is derived from existing data (e.g. based on old Wikipedia data or AI-generated data). Even when this approach is implemented successfully, new models are released regularly that are trained on more recent data, necessitating some analysis of train-test overlap with the previously novel test set. One modification of this approach is to add novel data to the test set at regular intervals, as with Livebench (White et al., 2024) or Image2Struct. [^3] See https://crfm.stanford.edu/helm/image2struct/latest/ In addition to the financial cost of continually adding novel data, which may not be feasible for every domain or project, one challenge of this approach is that it is difficult to interpret longitudinal progress.
Canary strings
as introduced by BIG-bench (Luo et al., 2024), are another strategy to cope with train-test overlap. Here, tasks in a test set are marked by a unique string, called a canary string, allowing developers to filter out data that contain canary strings during training. If a model outputs a given canary string, it signals that there is likely train-test overlap with the associated test set. But canary strings are not implemented uniformly or consistently—tasks exist without canary strings, whether within the test set or in other instances across the internet, and canary strings can be easily filtered out of test sets. More often test sets are derived from other raw sources that do not contain the canary string. It is also possible that canary strings may be referenced independently of the tasks in test sets, producing potential false positives.
Our position:
To complement these above approaches, language model developers should report train-test overlap statistics or openly release their training data. A developer chooses the specific test sets it uses to evaluate its language model, and it can choose to report train-test overlap for those test sets (e.g. through a transparency report or a model card) using its preferred method for computing train-test overlap (Bommasani et al., 2024b). This is similar to norms in the field of statistics, where published results must be accompanied by confidence intervals, rather than arbitrary reporting criteria imposed by a third party. This approach would complement existing strategies: for instance, black-box methods and canary strings are powerful tools to sanity check train-test overlap statistics that a developer reports. Similarly, private or novel test sets can further sanity check results on existing public test sets, such as drawing attention to cases of significant divergence.
3 Language Models
To establish a broad understanding of the landscape, we comprehensively consider the train-test overlap practices of the flagship language model of 30 developers (01.ai, Adept, AI2, AI21 Labs, Aleph Alpha, Alibaba, Amazon, Anthropic, Apple, BigCode, Cohere, Databricks, DeepSeek, EleutherAI, Technology Innovation Institute, Google, IBM, Imbue, Inflection, Meta, Microsoft, Mistral, NVIDIA, OpenAI, Reka AI, Snowflake, Stability AI, Together AI, Writer, and xAI). We assembled this list by considering all models on the HELM MMLU leaderboard [^4] See https://crfm.stanford.edu/helm/mmlu/latest/#/leaderboard and additional models that we selected for impact and relevance based on Ecosystem Graphs (Bommasani et al., 2023b).
Next, we selected the latest flagship model for which the developer had published benchmarks results. This is because we emphasize that a developer should disclose information about train-test overlap on the subset of benchmarks that the developer publishes rather than a pre-defined list of benchmarks decided by another party. For some developers, this meant selecting an older flagship model as there are as of yet no published results on the newer model. We chose to exclude developers that have not published results on public language benchmarks such as Baidu. We consider only models with results released before September 1, 2024 (the date we provided as a deadline to model developers to share additional train-test overlap information) and accordingly reported on Qwen2 rather than Qwen2.5, GPT-4 rather than GPT-4o  [^5] We note that GPT-4o system card was published before this date, but there is no new GPT-4o paper available yet so we chose to focus on GPT-4 instead., and OLMo rather than OLMoE.
4 Results
4.1 Documenting Current Practices
We followed a standardized procedure in order to document current practices regarding reporting of train-test overlap statistics. For each developer-model pair, we followed the following process to collate the developer’s current practices with respect to reporting train-test overlap:
  1. 1.
    We identified papers, technical reports, and company websites that were potential sources of information on train-test overlap.
  2. 2.
    We queried and identified any data the developer has published regarding the model’s results on public benchmarks. We documented each public benchmark on which the developer reports results for the model.
  3. 3.
    We queried each document that includes results on public benchmarks for information on train-test overlap. In addition to reading the document, we queried for terms including “contamination”, “overlap”, and “gram”, then manually inspected the occurrence to determine whether any train-test overlap data was released.
4.2 Scoring Criteria
We assign each developer a binary score of 1 or 0 to indicate whether the developer has provided sufficient information to contextualize train-test overlap for its flagship model. In this work we do not evaluate the specific methodology that each developer choses to employ to estimate train-test overlap, as these methodologies are inconsistent, opaque, and often not comprehensive. Instead, we identify whether a developer meets some minimum threshold with respect to publicly reporting some meaningful information about train-test overlap.
In assigning scores to developers, we consider the following criteria:
  1. 1.
    Is the training data publicly available?
  2. 2.
    Is train-test overlap reported on public benchmarks for which the model’s results are reported? That is, for each test set, we want a number that measures overlap. Note that this can be an implicit 0 for those who prefilter their training data.
    1. (a)
      Is train-test overlap reported with sufficient specificity to be meaningful?
    2. (b)
      Is there a clear description of the method the developer used to compute train-test overlap?
If none of these criteria are met, then the developer scores 0. If the training data is publicly available, the developer scores 1 as third parties can directly compute train-test overlap statistics for any public test set of interest. If the training data is not publicly available, but train-test overlap is reported with sufficient specificity and a clear description of the method, the developer scores 1.
For each developer that scored 0, we reached out to the developer to provide them an opportunity to engage with or rebut the score. Each of these developers was given the opportunity to point to any relevant information that our analysis was missing, or provide additional information that would be publicized.
4.3 Scores
Model Developer Score Explanation
Pythia EleutherAI 1 Open training data (Biderman et al., 2023)
OLMo AI2 1 Open training data (Groeneveld et al., 2024)
RedPajama-INCITE 7B Together AI 1 Open training data (Computer, 2024)
StarCoder 2 BigCode 1 Open training data (Lozhkov et al., 2024)
Palmyra X V3 Writer 1 Published analysis and code (Writer, 2024)
GPT-4 OpenAI 1 Published analysis (OpenAI et al., 2024)
Llama 3.1 Meta 1 Published analysis (Dubey et al., 2024)
Qwen2 Alibaba 1 Published analysis (Yang et al., 2024)
Apple Intelligence Apple 1 Published prefiltering (Gunter et al., 2024)
Gemini 1.5 Pro Google 0 Insufficient methodological details (Team et al., 2024a)
Arctic Snowflake 0 No analysis (Snowflake, 2024)
Claude 3.5 Sonnet Anthropic 0 No analysis (Anthropic, 2024)
Command R Cohere 0 No analysis (Cohere, 2024)
Core Reka AI 0 No analysis (Team et al., 2024b)
DBRX Databricks 0 No analysis (Databricks, 2024)
DeepSeek DeepSeek 0 No analysis (DeepSeek-AI et al., 2024)
Falcon TII 0 No analysis (Almazrouei et al., 2023)
Fuyu-Heavy Adept 0 No analysis (Adept, 2024)
Granite IBM 0 No analysis (Mishra et al., 2024)
Grok-2 xAI 0 No analysis (x.ai, 2024)
Imbue 70B Imbue 0 No analysis (Imbue, 2024)
Inflection-2.5 Inflection 0 No analysis (AI, 2024a)
Jamba-1.5 AI21 Labs 0 No analysis (AI21, 2024)
Luminous Supreme Aleph Alpha 0 No analysis (Alpha, 2024)
Mixtral Large 2 Mistral 0 No analysis (AI, 2024b)
Nemotron-4-340B-Instruct NVIDIA 0 No analysis (NVIDIA, 2024)
Phi 3 Microsoft 0 No analysis (Abdin et al., 2024)
Stable LM 2 Stability AI 0 No analysis (AI, 2024c)
Titan Text Express Amazon 0 No analysis (Amazon, 2024)
Yi-34B 01.ai 0 No analysis (AI et al., 2024)
Table 1: Models and scores. This table displays the score of 30 developers on our metric for train-test overlap transparency—a developer scores 1 if it releases sufficient information to contextualize for its flagship language model, and 0 otherwise. From left to right, the table includes: a list of flagship language models, a list of major model developers, the model developers’ scores, an abbreviated explanation for why the developer received that score.
Here we document the train-test overlap practices of 30 models with published results on public test sets. Of these, 9 models have published sufficient data for the community to contextualize train-test overlap: 4 models (OLMo—AI2, GPT-NeoX—EleutherAI, RedPajama INCITE—Together, StarCoder 2—BigCode/HuggingFace/ServiceNow) models have released training data under open-source licenses, which researchers can inspect for train-test overlap and 5 models (GPT-4—OpenAI, Llama 3.1—Meta, Qwen2—Alibaba, Palmyra—Writer, Apple Intelligence—Apple) have published their methodology and statistics for train-test overlap. For developers that do not openly release their training data, we provide additional explanation below as to why their transparency regarding train-test overlap is meaningful; for quantification of the degree of train-test overlap for each of these developers, see their associated technical reports.
OpenAI
reports its train-test overlap analysis in the GPT-4 Technical Report (OpenAI et al., 2024). OpenAI reports results for GPT-4 on 8 public test sets and shared train-test overlap analysis for 6 of these test sets. OpenAI et al. (2024) states “We measure cross-contamination between our evaluation dataset and the pre-training data using substring match. Both evaluation and training data are processed by removing all spaces and symbols, keeping only characters (including numbers). For each evaluation example, we randomly select three substrings of 50 characters (or use the entire example if it’s less than 50 characters). A match is identified if any of the three sampled evaluation substrings is a substring of the processed training example. This yields a list of contaminated examples. We discard these and rerun to get uncontaminated scores.”
Meta
reports its train-test overlap analysis in the Llama 3 Technical Report (Dubey et al., 2024, Section 5.1.4). Dubey et al. (2024) write: “Singh et al. (2024) propose to select contamination detection methods empirically, based on which method results in the largest difference between the ‘clean’ part of the dataset and the entire dataset, which they call estimated performance gain. For all our evaluation datasets, we score examples based on 8-gram overlap, a method that was found by Singh et al. (2024) to be accurate for many datasets. We consider an example of a dataset D to be contaminated if a ratio TD of its tokens are part of an 8-gram occurring at least once in the pre-training corpus. We select TD separately for each dataset, based on which value shows the maximal significant estimated performance gain across the three model sizes.” Meta reports train-test overlap for Llama 3.1 models on AGIEval, BIG-Bench Hard, BoolQ, CommonSenseQA, GSM8K, HellaSwag, MATH, NaturalQuestions, OpenBookQA, PiQA, QuaC, SiQA, SQuAD, Winogrande, and WorldSense.”
Writer
released train-test overlap statistics for Palmyra X after receiving a request from the authors. Writer ran a script [^6] This script is publicly available at https://github.com/stanford-crfm/data-overlap—we encourage other developers to run it over their training sets and publicly report the results. provided by the authors over its pretraining and instruction-tuning data to evaluate train-test overlap via an n-gram analysis; the results of this analysis are accessible at https://drive.google.com/file/d/1-3_GtZTyIbE1B5XWGGWtLRy5OwVtG0EK/view?usp=sharing. Writer publishes its results on HELM Lite (Writer, 2024), which includes 9 public test sets, and Writer reported train-test overlap on each of the public test sets included in HELM as well as others. Writer found some degree of train-test overlap for 13 of the 27 test sets on which it ran the script (APPS, CivilComments, CNN/Daily Mail, EntityMatching, HumanEval, ICE, LegalSupport, MATH, NarrativeQA, RAFT, The Pile, WikiFact, XSum).
Alibaba
released train-test overlap statistics for Qwen2 via an update to its technical report after receiving a request from the authors (Yang et al., 2024, Section 5.2.6). Yang et al. (2024) conducted an analysis of Qwen2’s training set following OpenAI’s approach, writing that in addition to n-gram matching “we also applied another constraint based on the longest common subsequence (LCS). Specifically, we first remove all symbols and punctuation from both the test and training sequences and perform tokenization. For a training sequence st, we remove it if there is a test sequence se such that $\lvert\text{LCS}(st,se)\rvert\geq 13$ and $\lvert\text{LCS}(st,se)\rvert\geq 0.6\times\min\left(\lvert st\rvert,\lvert se\rvert\right)$. To assess the potential effects of leaking data on the test performance, we follow OpenAI (2023) to construct a strict non-contaminated test set to check if there is a significant performance degradation after strict decontamination. Specifically, we construct the non-contaminated test set by excluding any sample which has 13-gram overlap with the pre-training or the post-training data (without constraint on LCS), and then compute the corresponding metric on the test set.” Alibaba reports results for Qwen2-72B on 14 public test sets (MMLU, GPQA, TheoremQA, HumanEval, MBPP, MultiPL-E, IFEval, LiveCodeBench v1, GSM8K, MATH, MT-Bench, MixEval, ArenaHard, and AlignBench) and reported train-test overlap on 8 of the public test sets (MMLU, GPQA, HumanEval, MBPP, MultiPL-E, GSM8K, MATH, and IFEval).
Apple
reports train-test overlap statistics for Apple Intelligence on 24 benchmarks (MMLU, GSM8K, HellaSwag, WinoGrande, NarrativeQA, Natural Questions, OpenBookQA, MATH_CoT, LegalBench, MedQA, WMT-2014, IFEval, AlpacaEval, ArenaHard, Berkeley Functional Calling, arc_challenge, arc_easy, lambada, piqa, sciq, triviaQA, webqs, HumanEval, MultiPLE-Swift); of these, Apple prefiltered its training data against 12 (MMLU, GSM8K, HellaSwag, WinoGrande, OpenBookQA, arc_challenge, arc_easy, lambada, piqa, sciq, triviaQA, webqs), filtering documents upon 4-13 gram collisions unless the n-gram reaches a “common-usage” threshold of 1000 (Gunter et al., 2024). Apple provided specificity about the benchmarks for which its training data was prefiltered after receiving a request from the authors.
5 Discussion
Overall, while train-test overlap is a fundamental to interpreting evaluation results, there is still significant limitations in the measurement methodology, beyond data access challenges and developer responsibility. As described above, direct string comparison is the most common way to quantify train-test overlap. This method has slight variations, but typically involves detecting substring matches between training and test data. N-gram matching is commonly employed (Yang et al., 2024; Dubey et al., 2024; Brown et al., 2020; Chowdhery et al., 2022), where documents are tokenized and then compared, though OpenAI compares characters rather than tokens in its analysis for GPT-4 (OpenAI, 2023). There are important design decisions developers make in employing n-gram strategies, such as what to set as N, whether to allow fuzzy matching or skipgrams (Dubey et al., 2024), and whether to filter based only a threshold of matches. This lack of uniformity in measurement approaches can make it challenging to directly compare train-test overlap analyses.
Benchmark Overlap Type Example
APPS Phrase (Question) Input—– The first line of the input contains a single integer n (1 $\leq$ n $\leq$ 100 000)
Phrase (Sequence) [’A’, ’B’, ’C’, ’D’, ’E’, ’F’, ’G’, ’H’, ’I’, ’J’, ’K’, ’L’, ’M’, ’N’, ’O’, ’P’, ’Q’, ’R’, ’S’, ’T’, ’U’, ’V’, ’W’, ’X’, ’Y’, ’Z’]
CivilComments Quote (News) Crews removed Baltimore’s Confederate statues early Wednesday, days after the deadly unrest in Charlottesville
EntityDataImputation Multi-Token Identifier AF-S DX Nikkor 18-55mm f/3.5-5.6G VR Lens
HumanEval Phrase (Math) prime number, or a prime, is a natural number greater than 1 that
Phrase (Sequence) ’zero’: 0, ’one’: 1, ’two’: 2, ’three’: 3, ’four’: 4, ’five’: 5, ’six’: 6, ’seven’: 7, ’eight’: 8, ’nine’: 9
ICE Phrase (Legal) with such intention, and under such circumstances that if by that act, you had caused the death of
LSAT Phrase (Question) Which one of the following could be a complete and accurate list of the
MATH Question Leakage 2-4+6-8+10-12+143-6+9-12+15-18+21
Phrase (Question) where $B$ is the area of the base and $h$ is the height
MMLU Question Leakage If K is a normal subgroup of H and H is a normal subgroup of G, then K is a normal subgroup of G.
NarrativeQA Quote (Book) M. Emanuel was away three years. Reader, they were the three happiest years of my life
RAFT Quote (News) the French language is very difficult but we try hard. If we come every day, maybe
RealToxicityPrompts Quote (News) When it comes to dashing the hopes of thousands of college-bound African
ThePile Phrase (Speech) If you have any questions, please feel free to give me a call.
Multi-Token Identifier the U.S. District Court for the Northern District of New York
Phrase (Sequence) Mime-Version: 1.0 Content-type: text/plain; charset=us-ascii Content-Disposition: inline
TwitterAAE Quote (Song) to be this way. Where did we go wrong we both made mistakes we gotta carry on
Summarization Quote (News) There are still many questions that the families of the 96 have and we believe that these people may be able to provide answers to some of those questions
Wikifact Quote (Wikipedia) swimming at the 1896 Summer Olympics – men’s sailors 100 metre freestyle
Table 2: Overlap Types and Examples on EleutherAI’s The Pile. For various test sets with overlap on The Pile, we chose contiguous overlapping n-gram sequences to illustrate the different types of overlap. Quotes refer to n-grams that exist in news, law, books, songs, or Wikipedia; Phrases refer to sequence of tokens that seem commonly grouped together; Multi-Token Identifier refers to a logical entity that is split into multiple tokens; and Question Leakage refers to an n-gram that may indicate that the core component of a question is overlapping. This was computed via https://github.com/stanford-crfm/data-overlap.
There are a number of limitations to this class of approaches. One limitation is that n-grams are coarse and do not capture the differences in types of overlap between different test sets. For instance, CivilComments is derived from news sites, so overlap between training data and CivilComments is likely due to news articles that appear in both the training and test data (Duchene et al., 2023). In contrast, MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) are in question-answer format, so overlap could stem from leakage of questions or answers or repetition of common phrases in questions and answers. We categorize the overlap types for different scenarios for The Pile (Gao et al., 2020) in 2. Overlap types can be broadly categorized into question leakage; quotes from news, laws, books, and songs; common phrases; and multi-token identifiers. Question leakage is the canonical concern for training data: if a model trains on the questions in the test set, it can achieve high results that fail to generalize well to new questions. However, these other overlap types can also be informative; for instance, simply matching the LSAT question stem “Which one of the following could be a complete and accurate list of the” suggests that the training data likely contains LSAT questions or similar questions. Indeed, train-test overlap is a construct that captures the relation between train and test data, and the different type of overlap add complexity that make it difficult to capture in a single statistic. Future work could explore these various overlap types in more detail and devise more granular metrics of measurement.
Another limitation is that n-gram analysis fails to catch many classes of train-test overlap that may be relevant, such as translations, summaries, or paraphrases of the text (Lee et al., 2024). Yang et al. (2023) demonstrate that there can be significant train-test overlap even with OpenAI’s prefiltering methods. Prior work has made progress on addressing this limitation, including by making use of embeddings or an LM-evaluator for a more semantic-based match (Dong et al., 2024; Jiang et al., 2024). These gaps demonstrate the need for further work on developing improved methods for estimating train-test overlap.
Indeed, in light of these limitations, we note that we are fundamentally interested in measuring the amount of generalization, rather than direct string matches or any specific approaches. This could extend to train-test overlap at the task or domain level. Additionally, it highlights that unlike the common perception, train-test overlap is not necessarily a negative (in part why we choose this term as opposed to “contamination”). Instead, it is helpful to guide understanding and help contextualize results.
Additionally, there are complexities in determining what qualifies as the training set, as there are often multiple stages of training and datasets, including pretraining, fine-tuning, and safety alignment among others (Yang et al., 2024; Dubey et al., 2024; Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023). This is often not captured in developers’ public reports (Bommasani et al., 2024a), and it may not be well captured internally either. Precision about the training set is important, though it is beyond the scope of this paper.
This paper does not assert a position on which method is best, and acknowledges that there is substantial research remaining to investigate better methods of computing train-test overlap. Nevertheless, the limitations of black-box approaches are far greater than those of white-box approaches. Just as the Foundation Model Transparency Index has helped improve the transparency of foundation model developers (Bommasani et al., 2023a; 2024a), our hope is that an increase in the number of model developers that report train-test overlap will produce better methods of measurement and help standardize reporting such that developers’ transparency on train-test overlap improves.
6 Conclusion
In this work, we highlight the need to improve transparency of train-test overlap. Our position is that any language model developer that publishes results on public test sets should release its training data and/or publish accompanying train-test overlap so that the community can interpret the results. We discuss various strategies to address train-test overlap, and how our position complements these efforts. We document the train-test overlap practices of 30 models with published results on public test sets. Of these, 9 models have published sufficient data for the community to contextualize train-test overlap. Finally, we discuss limitations with current approaches to quantifying train-test overlap, while emphasizing that current methods still have value. Instead, we suggest that as the AI community increasingly becomes aware of train-test overlap we can continue to improve upon and align on methodology for measuring and reducing train-test overlap.